OpenStudy (anonymous):

a)If log3x =2 then x=? b)If log6x=3 then x=?

OpenStudy (nenadmatematika):

a) x=100/3 b) x=500/3

OpenStudy (mertsj):

\[a) 10^{2}=3x\]

OpenStudy (mertsj):

Wait. Is 3 the base?

OpenStudy (anonymous):

nope

OpenStudy (mertsj):

Ok the. x = 100/3

OpenStudy (mertsj):

b) \[10^{3}=6x\]

OpenStudy (mertsj):

x=500/3

OpenStudy (anonymous):